In a geometric progression coasting of positive terms each term equals to thw sum of the next two terms then the common ratio is?

In a geometric progression coasting of positive terms each term equals to thw sum of the next two terms then the common ratio is?

Grade:10

1 Answers

Nayak Roshan
7 years ago
Hi Rural, 
Given a geometric progression, let the GP be a,ar, ar2,ar3,.....,arn.
Then the condition given here is a=ar+ar2 -(1)AND ar=ar2+ar(2l
Substuting 2 in one we get a = 2ar2+ar3-(3)
Dividing 3 and 2
a/ar=2ar2+ar3/ar2+ar3,
1/r=2+r/1+r,
1+r=2r+r2,
r2+r-1=0,
SOLVE THIS EQUATION AND THEN YOU WILL BE WITH YOUR ANSWER....
HOPE THIS HELPED YOU.....